# Simulating methylamine using symmetry adapted qubit-excitation-based variational quantum eigensolver

### v2

@@ Abstract
Understanding the capabilities of quantum computer devices and estimating the required resources to solve realistic tasks are among the most important problems towards achieving useful quantum computational advantage. In this work, we analyze the resources that are needed to simulate certain molecules on a medium-scale quantum computer with the use of the variational quantum eigensolver (VQE) approach. As the conventional realization of the VQE approach significant amount of resources for simulation, we propose and analyze optimization techniques based on molecular point group symmetries (symmetry adaption) and compact excitation circuits (qubit-excitation-based). These optimizations allows significant reduction of the essential computational capabilities yet ensuring the convergence to the required energies. We first apply this approach for small molecules, such as LiH and BeH2, to evaluate their compatibility, accuracy, and potential applicability to larger problems. Then we demonstrate that instead of 600,000 two-qubit operations (in the STO-3G basis using a naive version of the Unitary Coupled Cluster ansatz), we are able to simulate methylamine molecules with 26 qubits and about 12,000 of two-qubit gates using our method. We accomplish our analysis by estimating required resources for a formic acid molecule, whose simulation requires about 15,000 of two-qubit gates. Thus, the proposed combination of certain optimization methods can reduce the number of two-qubit operations by several orders of magnitude. Although, we present alternative approaches that are of interest in the context of the further optimization in the number of two-qubit operation, we note that these approaches do not perform well enough in terms of the convergence to required energies. While these challenges persist, our resource analysis represents a valuable step towards the practical use of quantum computers and the development of better methods for optimizing computing resources.
@@ I Introduction
A problem of calculating energies and electronic structures of molecular ground and excited states is of primary importance for various industry-relevant problems, such as finding catalysts for chemical reactions and materials design Helgaker et al. (2000). However, the computational complexity of exact quantum chemistry simulations for many relevant situations grows exponentially with the system size Kempe et al. (2006). This problem has stimulated the development of classical approaches that capture relevant chemical properties van Gunsteren and Oostenbrink (2024), yet used simplifications may hinder some underlying electronic correlation effects that may play important role for various cases.
To overcome these limitations and address the problem of growing computational demands, the quantum computing paradigm present a promising solution Lloyd (1996); Cao et al. (2019); McArdle et al. (2020). By leveraging the principles of quantum physics, quantum computing devices offer the potential to efficiently solve complex quantum-chemical problems that are intractable for classical computers Cao et al. (2019); McArdle et al. (2020).
@@
In this work, we analyze the resources that are needed to simulate certain molecules on a medium-scale quantum computer with the use of the VQE approach taking into account possible optimization. Specifically, we combine the use of molecular point group symmetries (symmetry adaption) and compact excitation circuits (qubit-excitation-based). As we demonstrate, these optimizations allows significant reduction of the essential computational capabilities yet ensuring the convergence to the required energies. First we use this approach for small molecules, such as LiH and BeH2, to evaluate their compatibility, accuracy, and potential applicability to larger problems. We then simulate methylamine molecules with 30 qubits and about 12,000 of two-qubit gates using our method, which is the significant gain the terms of required resources compared to the standard approaches. We accomplish our analysis by estimating required resources for a formic acid molecule, whose simulation requires about 15,000 of two-qubit gates. The inter-nuclear coordinates used in this study have been sourced from the CCCBDB database III (2022) and optimized at the CCSD(T) level of theory with the Def2TZVPP basis set. Our results demonstrate that these optimization techniques performed well and reduce the circuit depth and the number of two-qubit gates by an order of magnitude making such simulations more feasible using near-term quantum computers.
@@
Our paper is organized as follows. We formulate the problem in Sec. II. In Sec. III, we formulate the VQE approach for solving quantum-chemical problems. In Sec. IV, we analyze the required quantum resourced for simulating methylamine and formic acid molecules using the VQE approach. We take into account possible optimization techniques that can be used to reduce the amount of the required results for VQE-based quantum chemical simulations. We summarize our results and conclude in Sec. V.
II Electronic wave function
In the most general case, the electronic wave function $\Psi$ for a many-body system, such as an $N$-electron molecule, can be expressed as a superposition of multiple Slater determinants as follows:
$\Psi(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})=\sum_{I}d_{I}\mathcal{A}\Bigl{\{}\chi_{1}^{(I)}(\mathbf{x}_{1})\chi_{2}^{(I)}(\mathbf{x}_{2})\\ \cdots\chi_{N}^{(I)}(\mathbf{x}_{N})\Bigr{\}}.$ (1)
where $\mathbf{x}_{k}=(\mathbf{r}_{k},\omega_{k})$ includes spatial and spin coordinates, $\chi_{k}(\mathbf{x}_{k})$ is a single-electron wave function (spin-orbital), and $\mathcal{A}\{\dots\}$ denotes antisymmetrization (the Slater determinant). The coefficients $d_{I}$ of the determinants can represent certain variational or combinatorial coefficients. Each individual term in this sum corresponds to a particular electronic configuration of the molecular electronic system (ground state and various excited states).
This generalized representation is valid in the context of the Full Configuration Interaction (FCI) method. However, a full linear combination of determinants is rarely used in practical quantum chemical calculations because the complexity of FCI grows exponentially with the number of electrons and orbitals.
In the context of the VQE approach, the initial approximation often involves a single determinant corresponding to the ground state of the molecular system, which is equivalent to solving the system using the Hartree-Fock method. The wave function (statevector) in this case is then given by:
$|\Psi_{\mathrm{HF}}\rangle=\frac{1}{\sqrt{N!}}\begin{vmatrix}\chi_{1}(\mathbf{x}_{1})&\chi_{2}(\mathbf{x}_{1})&\cdots&\chi_{N}(\mathbf{x}_{1})\\ \chi_{1}(\mathbf{x}_{2})&\chi_{2}(\mathbf{x}_{2})&\cdots&\chi_{N}(\mathbf{x}_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \chi_{1}(\mathbf{x}_{N})&\chi_{2}(\mathbf{x}_{N})&\cdots&\chi_{N}(\mathbf{x}_{N})\end{vmatrix}.$ (2)
Each single-electron function (spin-orbital) $\chi_{k}(\mathbf{r}_{k},\omega_{k})$ within the determinant can be expressed as a product of two components, where:
  1. 1.
    $\psi_{k}(\mathbf{r}_{k})$ is the molecular orbital (MO), which describes the spatial distribution of electron density.
  2. 2.
    $\xi_{k}(\omega_{k})$ is one of the two spin functions (commonly denoted as $\alpha$ or $\beta$), accounting for the possible projections of the electron’s spin angular momentum.
Thus, we can write the expression as follows:
$\chi_{k}(\mathbf{r_{k}},\omega_{k})=\psi_{k}(\mathbf{r}_{k})\,\xi_{k}(\omega_{k}).$ (3)
Next, an approximation for the spatial part (MO) of the single-electron spin-orbital using the Linear Combination of Atomic Orbitals (LCAO) method can be used:
$\psi_{k}(\mathbf{r}_{k})=\sum_{\mu}C_{\mu k}\,\phi_{\mu}(\mathbf{r}_{k}),$ (4)
where $\phi_{\mu}(\mathbf{r}_{k})$ are atomic orbitals (AOs), and $C_{\mu j}$ are coefficients determined, for example, from the Roothaan equations (Hartree-Fock method). Further representation of AOs depends on the chosen basis set.
The orbitals derived from such approximations serve as input parameters for quantum computational algorithms, encoding the electronic structure into a qubit representation. However, for quantum algorithms based on phase estimation, the number of qubits and quantum operations grows rapidly with system size, making them computationally infeasible for the current and foreseen hardware Reiher et al. (2017). To address this challenge, hybrid quantum-classical approaches have been introduced Cerezo et al. (2021a). One of the most widely adopted of these is the VQE, which leverages both quantum hardware and classical optimization and is presented in the section below.
III VQE formalism
@@
During the last decade, a large variety of modifications of the VQE approach have been proposed Cerezo et al. (2021a), however the majority of them is based on the Ritz variational principle. The Ritz principle involves choosing a certain parametrized combination of wave functions or state vectors i.e. the ansatz, and minimizing the functional defined on this parametrized combination. Choosing the right ansatz should take into account the specifics of the problem and its boundary conditions. In other words, the ansatz represents a linear combination of some known functions, which are parameterized by some unknown coefficients:
@@
To transform fermionic operators into qubit operators, several transformations, such as Jordan-Wigner, Parity, and Bravyi-Kitaev, can be used. Qubit operators or Pauli strings, refer to tensor products of Pauli matrices $X$, $Y$, $Z$ and an identity operator $I$ that act on qubits. After the transformation is realized, one obtain a qubit Hamiltonian in the form
$H=\sum_{i}\beta_{i}P_{i},$ (9)
where $P_{i}$ are Pauli strings. Each Pauli string can be then measured individually or in commutative groups using additional post-rotation gates, preceded by the same ansatz circuit for every Pauli string.
@@
$\hat{T}(\bm{\theta})=\hat{T}_{1}(\bm{\theta})+\hat{T}_{2}(\bm{\theta}).$ (11)
@@ IV VQE simulations
The number of qubits can be reduced by employing the frozen core approximation, where the inner 1s orbitals of carbon and oxygen in the formic acid molecule and carbon and nitrogen in the methylamine molecule are frozen. This step reduces the required number of qubits by 6 for CH2O2 and by 4 for CH3NH2. Additionally, by applying a technique known as qubit tapering Bravyi et al. (2017), which exploits the discrete Z2 symmetries of the Hamiltonian Pauli strings, we can further eliminate three additional qubits from our simulation for both molecules, resulting in final counts of 25 and 23 qubits.
@@
Table 1: Calculated atomic coordinates for methylamine (CH3NH2) and formic acid (CH2O2) in angstroms.
Utilizing the coordinates in Table 1 allows us to derive the qubit form of the molecular Hamiltonian (9) and ascertain the number of terms it contains. For molecular Hamiltonians it scales with the number of electrons as $\mathcal{O}(N^{4})$. This is a critical factor in determining the requisite number of quantum circuits to be measured at each iteration of VQE algorithm. Besides, we can adopt more sophisticated grouping methods for Pauli terms, such as qubit-wise and general groupings Gokhale et al. (2020). These methods leverage the commutative properties of Pauli terms to notably decrease the number of necessary measurements.
@@
The total circuit depth can be estimated in the similar way:
@@
where $\gamma_{S},\;\gamma_{D}$ contribution to the circuit depth from each single/double excitation. The coefficients from Eqs. (13) and (14) depend on the actual definition of excitation sub-circuits in the ansatz, the fermionic transformation used, and the number of qubits. Recent papers have presented more optimal ways of constructing such circuits with fewer CNOT gates Yordanov et al. (2020, 2021); Magoulas and Evangelista (2023).
@@
One could use so-called qubit excitations, which always have a fixed number of CNOT gates. However, it is not clear whether they can adequately account for the electronic correlation contribution, as they ignore the anticommutation of fermions by excluding Pauli Z products from the operator exponent.
Parameter CH3NH2 CH2O2
Ansatz Type UCCSD UpGSD UCCSDQs UpGSDQ UpGSDQs UCCSD UpGSD UCCSDQs UpGSDQ UpGSDQs
Circuit Depth 514379 14758 5766 473 247 563722 15026 7563 514 266
Qubits 23 23 26 26 26 25 25 28 28 28
Total Gates 631022 18526 11558 2822 1526 682524 19466 14610 3294 1638
Two-qubit gates 457800 12896 4130 1014 546 510426 13520 5221 1183 585
Pauli Terms 20908 20908 20908 20908 20908 30423 30423 30423 30423 30423
Circuits 4144 4144 4144 4144 4144 5103 5103 5103 5103 5103
Double Excitations 2394 78 314 78 42 2745 91 397 91 45
Tapering Yes Yes No No No Yes Yes No No No
Table 2: Estimated resources for the VQE computation of CH3NH2 and CH2O2 molecules. The number of qubits is calculated using the frozen core approximation and the qubit tapering procedure. The number of circuits is based on qubit-wise operator grouping. Parameters for the k-UpGSD, k-UpGSDQ, and k-UpGSDQs ansatzes are estimated with $k=1$. The best-performing method in terms of the trade-off between required resources and VQE convergence to the FCI energy is highlighted in bold.
A naive implementation of the UCCSD ansatz leads to excessive resource requirements since the circuit depth (14) is proportional to the number of two-qubit gates which scales as $\sim\mathcal{O}(N^{4})$. However, one can employ several techniques to slightly reduce these requirements. One of them is k-UpCCGSD (for brevity k-UpGSD) Lee et al. (2018); Burton et al. (2023) ansatz, where $k$ stands for the number of repetitions of UpCCGSG ansatz with the new set of variational parameters for each repetition.
The main difference from UCCSD is that it includes only two-body terms, which shift pairs of opposite-spin electrons from fully occupied to completely unoccupied spatial orbitals making sure there are no singly occupied states described. The estimated circuit parameters for this method are presented in Table 2. Its combination with the qubit excitations approach was also evaluated and is shown in the same table. However, the qubit tapering procedure was not applied to qubit excitations due to incompatibilities between these methods.
An alternative way to optimize the UCCSD ansatz is to filter out the excitation terms by their irreducible representations Cao et al. (2022). In other words, one should retain only those terms whose irreducible representations coincide with the irreducible representation of the target state It is presented in combination with the previous methods as k-UpGSDQs in Table 2. Moreover, the use of the Bravyi-Kitaev mapping could result in fewer two-qubit operations count due to lower Pauli-weight Seeley et al. (2012), i.e. the maximum number of non-identity operators in any Pauli string produced by the mapping, in some cases it can improve scaling of Eq. (13).
For the current resource estimations, an in-house quantum chemistry library was utilized, specifically developed for the efficient implementation of variational quantum algorithms Makushin et al. (2024). First, to make sure the chosen ansatz construction methods and optimization techniques work as expected, we tested them on two relatively simple molecules, LiH and BeH2, both with a small number of SO (STO-3G basis suggested for both molecules). For LiH, we consider freezing the inner 1s orbital, which resulted in 10 qubits in the non-tapered regime and 6 in the tapered regime. For BeH2, we initially have 14 qubits; the frozen core method reduced this to 12 qubits in the non-tapered regime and 7 qubits with tapering. We then calculate their ground state energies using the VQE and compared the results for different ansatz types. For both molecules, we use a noiseless statevector simulator in the combination with the BFGS algorithm for classical optimization.
@@
We present the results of VQE calculations for the LiH molecule in Fig. 1. The plot shows how the energy converged over iterations for various ansatz types, such as UCCGSD, UCCSD, 2-Up-GSD, UCCSDQ, and UCCSDQs. The dashed black line represents the exact FCI energy, which we used as a reference. The shaded area around it shows the chemical accuracy threshold of $\pm$0.0016 Hartree. Among all the ansatz types, UCCSDQs demonstrates the best performance: it achieves the lowest circuit depth (137) and two-qubit gate count (86), while still maintaining chemical accuracy.
@@
In Fig. 2, we present the same analysis for the BeH2 molecule. Again, one can see how each ansatz converged to the ground state energy. As in the case of the LiH simulation, UCCSDQs proves to be the most resource-efficient, as it requires the least resources (circuit depth of 279 and two-qubit gate count of 190), while staying within the chemical accuracy range The energy curves initially drop sharply and then then level off as they approach the exact FCI value, highlighting the efficiency of these methods for BeH2.
@@
Overall, these results show that the methods that we use perform well for calculating ground state energies of small molecules like LiH and BeH2222. They also highlight how different ansatz types can affect the balance between accuracy and resource needs. While these methods perform great for smaller systems, scaling them up to larger molecules is a whole different challenge.
Indeed, as shown in Table 2, which gives resource estimates for more complex molecules, such as methylamine and formic acid, the computational demands become significant. Even with the STO-3G minimal basis set, the required circuit depth and number of qubits for these molecules are way beyond what current quantum hardware can handle. Alternative methods for constructing the ansatz, such as adaptive ansatzes (ADAPT-VQE) Grimsley et al. (2019); Sapova and Fedorov (2022) or evolutionary algorithms Chivilikhin et al. (2020), can reduce the computational resources, for example by shortening the circuit depth or reducing the total number of variational parameters. However, these methods can incur a higher cost in terms of the number of circuit measurements, because each step in the adaptive or evolutionary procedure often requires extra loops and operator gradient evaluations.
@@
Hence, if we let $K$ be the total number of ADAPT-VQE iterations and $P$ be the operator pool size, the naive measurement cost can scale as
$\mathcal{O}\bigl{(}K\times P\bigr{)}$ (15)
in gradient evaluations, where each gradient evaluation might involve an inner VQE with potentially many measurements. If each VQE step needs $M_{\text{VQE}}$ measurements for its Hamiltonian estimation, and each gradient check does $m$ param-shift or finite-difference evaluations, we might end up with
$\mathcal{O}\bigl{(}K\times P\times m\times M_{\text{VQE}}\bigr{)}$ (16)
as a rough total measurement scaling. In terms of the number of qubits, each of these multiples is proportional to $N_{q}$ raised to some power $\alpha$, so the total number of measurements scales polynomially, $\mathcal{O}\bigl{(}N_{q}^{\alpha}\bigr{)}$.

### v3

@@ Abstract
Understanding the capabilities of quantum computer devices and estimating the required resources to solve realistic tasks are among the most important problems towards achieving useful quantum computational advantage. As the conventional realization of the variational quantum eigensolver (VQE) approach requires a significant amount of resources, we propose and analyze optimization strategies that combine various methods, including molecular point group symmetries (symmetry adaptation), compact excitation circuits (qubit-excitation-based), different types of excitation sets, and qubit tapering. These strategies allow for a significant reduction in computational requirements while ensuring convergence to the correct energies. First, we apply these combinations to small molecules, such as LiH and BeH2, to evaluate their compatibility, accuracy, and potential applicability to larger problems. We then simulate the methylamine molecule within its restricted active space using the best-performing optimization strategies. Finally, we complete our analysis by estimating the resources required for full active-space simulations of the methylamine and formic acid molecules. Our best-performing optimization strategy reduces the number of two-qubit operations for simulating the methylamine molecule from 600,000 (in the STO-3G basis with a naive Unitary Coupled Cluster ansatz) to approximately 12,000, using 26 qubits. Thus, the proposed combination of optimization methods can reduce the number of two-qubit operations by nearly two orders of magnitude. Although, we present alternative approaches that are of interest in the context of the further optimization in the number of two-qubit operation, we note that these approaches do not perform well enough in terms of the convergence to required energies. While these challenges persist, our resource analysis represents a valuable step towards the practical use of quantum computers and the development of better methods for optimizing computing resources.
@@ I Introduction
The problem of calculating energies and electronic structures of molecular ground and excited states is of primary importance for various industry-relevant problems, such as finding catalysts for chemical reactions and materials design Helgaker et al. (2000). However, the computational complexity of exact quantum chemistry simulations for many relevant situations grows exponentially with the system size Kempe et al. (2006). This problem has stimulated the development of classical approaches that capture relevant chemical properties van Gunsteren and Oostenbrink (2024), yet used simplifications may hinder some underlying electronic correlation effects that may play important role for various cases.
To overcome these limitations and address the problem of growing computational demands, the quantum computing paradigm presents a promising solution Lloyd (1996); Cao et al. (2019); McArdle et al. (2020). By leveraging the principles of quantum physics, quantum computing devices offer the potential to efficiently solve complex quantum-chemical problems that are intractable for classical computers Cao et al. (2019); McArdle et al. (2020).
@@
In this work, we analyze the resources that are required to simulate certain molecules on a medium-scale quantum computer with the use of the VQE approach, taking into account possible optimization. Specifically, we combine the use of molecular point group symmetries (symmetry adaptation), compact excitation circuits (qubit-excitation-based), several types of excitation sets, and the qubit tapering method. As we demonstrate, these optimizations allow for a significant reduction in computational requirements while ensuring the convergence to the required energies. First, we apply these combinations to small molecules, such as LiH and BeH2, to evaluate their compatibility, accuracy, and potential applicability to larger problems. We then simulate the methylamine molecule within its restricted active space using the best-performing optimization strategies. Finally, we complete our analysis by estimating the resources required for full active-space simulations of the methylamine and formic acid molecules, which, using our best optimization strategy, would require around 10,000–15,000 two-qubit gates. Both the two-qubit gate count and circuit depth are improved by nearly two orders of magnitude compared to the naive approach. Our results demonstrate that these optimization strategies perform well and can make such simulations more feasible for near-term quantum computers.
@@
The paper is organized as follows. In Sec. II, we formulate the VQE approach for solving quantum-chemical problems. In Sec. III, we analyze the required quantum resourced for simulating methylamine and formic acid molecules using the VQE approach. We take into account possible optimization techniques that can be used to reduce the amount of the required results for VQE-based quantum chemical simulations. We summarize our results in Sec.IV and present our conclusions in Sec.V.
II VQE formalism
@@
During the last decade, a large variety of modifications of the VQE approach have been proposed Cerezo et al. (2021a), the majority of them is based on the Ritz variational principle. The Ritz principle involves choosing a certain parametrized combination of wave functions or state vectors i.e. the ansatz, and minimizing the functional defined on this parametrized combination. Choosing the right ansatz should take into account the specifics of the problem and its boundary conditions. In other words, the ansatz represents a linear combination of some known functions, which are parameterized by some unknown coefficients:
@@
To transform fermionic operators into qubit operators, several mappings, such as Jordan-Wigner, Parity, and Bravyi-Kitaev, can be used. After the transformation is realized, one obtain a qubit Hamiltonian in the form:
$H=\sum_{i}\beta_{i}P_{i},\quad\text{with}\quad P_{i}=\bigotimes_{k=1}^{N}\sigma_{k}^{(i)},$ (5)
where each $\sigma_{k}^{(i)}\in\{I,X,Y,Z\}$ is a Pauli operator, $N$ is a number of qubits, and $P_{i}$ are Pauli strings. Each Pauli string can be then measured individually or in commutative groups using additional post-rotation gates, preceded by the same ansatz circuit for every Pauli string.
@@
$\hat{T}(\bm{\theta})=\hat{T}_{1}(\bm{\theta})+\hat{T}_{2}(\bm{\theta})=\sum_{i}\theta_{i}P_{i}.$ (7)
@@ III VQE simulations
The number of qubits can be reduced by employing the frozen core approximation, where the inner 1s orbitals of carbon and oxygen in the formic acid molecule and carbon and nitrogen in the methylamine molecule are frozen. This step reduces the required number of qubits by 6 for CH2O2 and by 4 for CH3NH2.
Additionally, by applying a technique known as qubit tapering Bravyi et al. (2017), one can exploit symmetries of the system to find a unitary transformation $U$, which acts on the Hamiltonian (5) such that:
$H^{\prime}=UHU^{\dagger}=\sum_{i}\beta^{\prime}_{i}P^{\prime}_{i},$ (8)
where $H^{\prime}$ has the same eigenvalues as $H$. After this transformation, there are operators acting trivially in each Pauli string $P^{\prime}_{i}$ of the transformed Hamiltonian. In other words, the Pauli strings have either the identity operator $I$ or the same Pauli operator (e.g., $X$) at specific qubit positions across all $H^{\prime}$ terms. These operators can be replaced by their eigenvalues, and the corresponding qubits can be effectively removed from the simulation. In general, if there are $m$ independent $\mathbb{Z}_{2}$ symmetries, the number of active qubits is reduced by $m$. In the case of formic acid and methylamine molecules, we can eliminate three additional qubits from our simulation for both molecules, resulting in final counts of 25 and 23 qubits, respectively.
@@
Table 1: Calculated atomic coordinates for methylamine (CH3NH2) and formic acid (CH2O2) in angstroms. The coordinates sourced from the CCCBDB database III (2022) and optimized at the CCSD(T) level of theory with the Def2TZVPP basis set.
Utilizing the coordinates in Table 1 allows us to derive the qubit form of the molecular Hamiltonian (5) and ascertain the number of terms it contains. For molecular Hamiltonians it scales with the number of electrons $N$ as $\mathcal{O}(N^{4})$. This plays a central role in defining the required number of quantum circuits to be measured at each iteration and represents one of the major bottlenecks in the VQE implementation for complex molecules. To partially overcome this problem one can adopt more sophisticated grouping methods for Pauli terms, such as qubit-wise and general groupings. These methods leverage the commutative properties of Pauli terms to notably decrease the number of necessary measurements, scaling as $\sim\mathcal{O}(N^{4}/30)$ for large Hamiltonians. We can say that two Pauli strings qubit-wise commute if, at each index, the corresponding two Pauli operators commute, for instance, $\{YY,YI,IY,II\}$. In the case of general commutativity, two Pauli strings commute if and only if they do not commute on an even number of indices, for example, $\{XX,YY,ZZ\}$ Gokhale et al. (2020). Such sets can be measured simultaneously with a single circuit, leveraging their commutativity.
@@
The total circuit depth can be estimated in a similar way:
@@
where $\gamma_{S},\;\gamma_{D}$ are the contributions to the circuit depth from single and double excitations, respectively. The coefficients from Eqs. (10) and (11) depend on the actual definition of excitation sub-circuits in the ansatz, the fermionic transformation used, and the number of qubits. Standard implementations of exponentials of Pauli strings (6-7) in quantum circuits typically rely on sequences of single-qubit rotations and cascades of CNOT gates arranged in a ladder-like structure. Recent papers have presented optimal ways of constructing such circuits with fewer CNOT gates Yordanov et al. (2020, 2021); Magoulas and Evangelista (2023). One approach is to use qubit excitations, which always require a fixed number of CNOT gates. However, it remains unclear whether they can adequately account for the electronic correlation contribution, as they ignore the anticommutation of fermions by excluding Pauli-Z products from the operator exponent.
Parameter CH3NH2 CH2O2
Ansatz Type UCCSD k-UpGSD UCCSDQs k-UpGSDQ UCCSD k-UpGSD UCCSDQs k-UpGSDQ
Circuit Depth 514379 14758 5766 473 563722 15026 7563 514
Qubits 23 23 26 26 25 25 28 28
Total Gates 631022 18526 11558 2822 682524 19466 14610 3294
Two-qubit gates 457800 12896 4130 1014 510426 13520 5221 1183
Pauli Terms 20908 20908 20908 20908 30423 30423 30423 30423
Circuits 4144 4144 4144 4144 5103 5103 5103 5103
Double Excitations 2394 78 314 78 2745 91 397 91
Tapering Yes Yes No No Yes Yes No No
Table 2: Estimated resources for the VQE computation of CH3NH2 and CH2O2 molecules. The number of qubits is calculated using the frozen core approximation and the qubit tapering procedure. The number of circuits is based on qubit-wise operator grouping. The parameters for the k-UpGSD and k-UpGSDQ ansatzes are estimated with $k=1$, To obtain estimates for higher values of $k$, the table values should be multiplied by the corresponding scaling factor. The best-performing method in terms of the trade-off between required resources and VQE convergence to the FCI energy is highlighted in bold.
Alternative methods for constructing the variational ansatz, such as adaptive ansatzes (ADAPT-VQE) Grimsley et al. (2019); Sapova and Fedorov (2022) or evolutionary algorithms Chivilikhin et al. (2020), can reduce the computational resources, for example by shortening the overall circuit depth (11) or reducing the total number of variational parameters. However, they can incur a higher cost in terms of the number of circuit measurements, because each step in the adaptive or evolutionary procedure often requires extra loops and operator gradient evaluations.
A naive implementation of the UCCSD ansatz leads to excessive resource requirements since the circuit depth (11) is proportional to the number of two-qubit gates which scales as $\mathcal{O}(N^{4})$. However, one can employ several techniques to reduce these requirements. One of them is k-UpCCGSD (for brevity k-UpGSD) Lee et al. (2018); Burton et al. (2023) ansatz,where $k$ represents the number of ansatz repetitions, each with a new set of variational parameters, $p$ refers to paired excitations, and $G$ denotes the inclusion of generalized excitations. The main difference from UCCSD is that it includes only two-body terms, which shift pairs of opposite-spin electrons from fully occupied to completely unoccupied spatial orbitals making sure there are no singly occupied states described. The ansatz circuit’s depth is reduced because the number of paired generalized excitations is significantly smaller than that of standard UCCSD excitations.
Another way to optimize the UCCSD ansatz is to filter out the excitation terms by their irreducible representations Cao et al. (2022). In other words, one should retain only those terms whose irreducible representations coincide with the irreducible representation of the initial state (Hartree-Fock determinant), in the case of doubly occupied molecular orbitals we can write:
$\forall\Gamma\left(e^{\hat{T}-\hat{T}^{\dagger}}|\psi_{HF}\rangle\right)=\Gamma(|\psi_{HF}\rangle):e^{\hat{T}-\hat{T}^{\dagger}}=1,$ (12)
here $\Gamma$ denotes the irreducible irreducible representation of the corresponding wavefunction, ensuring that only excitations preserving the symmetry of the reference state $|\psi_{HF}\rangle$ are valid in the UCCSD ansatz.
IV Results
We propose several optimization strategies that realize different compatible combinations of the aforementioned methods. Qubit excitations are designated as $Q$, and irreducible representation filtering as $s$, which are combined with UCCSD and k-UpGSD ansatzes to form UCCSDQ, UCCSDQs and k-UpGSDQ. Additionaly we make resource estimations for UCCSD and k-UpGSD ansatzes in combination with qubit tapering and compare them with other optimization strategies. In all estimates, we use qubit-wise grouping of Pauli strings and an in-house quantum chemistry library specifically developed for the efficient implementation of variational quantum algorithms. Makushin et al. (2024).
First, to make sure the chosen ansatz construction methods and optimization strategies work as expected, we test them on two relatively simple molecules, LiH and BeH2, both with a small number of SO (STO-3G basis suggested for both molecules). For LiH, we consider freezing the inner 1s orbital, which results in 10 qubits in the non-tapered regime and 6 in the tapered regime. For BeH2, we initially have 14 qubits; the frozen core method reduces this to 12 qubits in the non-tapered regime and 7 qubits with tapering. We then calculate their ground state energies using the VQE and compare the results for different ansatz types. For both molecules, we use a noiseless statevector simulator in the combination with the BFGS algorithm for classical optimization.
@@ IV Results
The results of the VQE calculations for the LiH molecule are presented in Fig. 1. The plot shows how the energy converges over iterations for various ansatz types, such as UCCSD (plus its generalized version UCCGSD), 2-UpGSD, 2-UpGSDQ, UCCSDQ, and UCCSDQs. The dashed black line represents the exact FCI energy, which we use as a reference. The shaded area around it shows the chemical accuracy threshold of $\pm$0.0016 Hartree. Among all the ansatz types, UCCSDQs demonstrates the best performance: it achieves the lowest circuit depth (137) and two-qubit gate count (86), while still maintaining chemical accuracy.
@@
In Fig. 2, we present the same analysis for the BeH2 molecule. Again, one can see how each ansatz converges to the ground state energy. The k-UpGSDQ ansatz requires 4 circuit repetitions to achieve convergence. As in the case of the LiH simulation, UCCSDQs proves to be the most resource-efficient, as it requires the least resources (circuit depth of 279 and two-qubit gate count of 190), while staying within the chemical accuracy range. The energy curves initially drop sharply and then level off as they approach the exact FCI value, highlighting the efficiency of these methods for BeH2.
@@
Overall, these results show that the optimization strategies we use perform well for calculating ground state energies of small molecules like LiH and BeH2. They also highlight how different ansatz types can affect the balance between accuracy and resource needs. While these strategies perform great for smaller systems, scaling them up to larger molecules is a whole different challenge. Indeed, as shown in Table 2, which gives resource estimates for more complex molecules, such as methylamine and formic acid, the computational demands become significant. Even with the STO-3G minimal basis set, the required circuit depth and number of qubits for these molecules are way beyond what current quantum hardware can handle.
@@
The use of more complex bases, that provide a better description of valence orbitals or account for polarization effects and electron correlations in molecules, such as 6-31G, 6-31G*, or cc-pVDZ, is associated with similar difficulties. All of them imply that significantly more qubits are required compared to the minimal basis case. For instance, simulating the H4 molecule using the cc-pVDZ basis would require 40 qubits Guo et al. (2023). For simulating the H2O molecule in the 6-31g basis, 50 qubits are needed. For methylamine and formic acid molecules the number of qubits exceed 60 in the case of complex bases. Recent studies Li et al. (2022); Shang et al. (2023) investigated various ways to enable the simulation of relatively large systems with small computational resources. A commonly used approach in these works is the HOMO-LUMO approximation, which restricts the active space to a subset of orbitals of a molecule. This method results in a more compact qubit representation of the molecular system, making simulations more feasible. However, this reduction in active space introduces additional approximation errors, which can impact accuracy.
Similarly, we demonstrate the results of simulations for the Methylamine molecule with a restricted number of spin orbitals. We take an active space of 6 electrons and 6 spin orbitals from the full 15 spin-orbital space. First, we will show the energy difference for such a choice. In Fig. 3, one can observe different configurations of the methylamine active space and their corresponding energies. This choice of active space reduces the number of qubits required to simulate the molecular system. However, as the number of spin orbitals decreases, the calculated energy becomes closer to the Hartree-Fock result. Chemical accuracy is achieved only when the core 1s orbitals are frozen. For the (6e, 6o) active space configuration of methylamine we use the UCCSDQ and UCCSDQs ansatzes in VQE calculation and the same minimal basis STO-3G. As it shown in Fig 4 the energy converges to the FCI value for both ansatzes. This result demonstrates that for the 12-qubit methylamine active space, the VQE algorithm, combined with the aforementioned absatzes, can achieve chemical accuracy within a small number of iterations. Other types of ansatzes are not tested due to higher computational cost.